Please help me to solve the question #5 with an explanation:

Three charges (q1 = 5.8 pC q2 =-5.9 pC and q3-3.4pC are located at the vertices of an equilateral triangle with side d 9.3 cm as shown. yt 43 91 1) What is F3,x the value of the x-component of the net force on q3? 20.70 Submit 2) What is F3,y, the value of the y-component of the net force on q3? 0.3064 Submit

3) q3 2 A charge q4 3.4 HC is now added as shown What is F2,x, the x-component of the new net force on q2? -56.4828 NI Submit 4) What is F2,y, the y-component of the new net force on q2? Submit 5) What is F1,x%, the x-component of the new net force on q1? N Submit

Thank you!
Answer #1

given

q1 = 5.8 uC

q2 = - 5.9 uC

q3 = 3.4 uC

d = 9.3 cm = 0.093 m

cos \theta = d/2 / d

\theta = 60o

a )

F3x = k q3 ( q1 + q2 ) cos \theta/ d2

= 9 x 109 x 3.4 x 10-6 x ( 5.8 + 5.9 ) x 10-6 x cos60 / 0.0932

F3x = 20.69 N

b )

F3y = k q3 ( q1 - q2 ) sin \theta/ d2

= 9 x 109 x 3.4 x 10-6 x ( 5.8 -0.0 5.9 ) x 10-6 x sin60 / 0.0932

F3y = - 0.306 N

c )

F2x = - ( k q2 ( q1 + q2 + q3 + q4 ) cos \theta/ d2

= - 9 x 109 x 5.9 x 10-6 x cos60 x ( 5.8 + 5.9 + 3.4 + 5.9 ) x 10-6 / 0.0932

F2x = - 64.46 N

d )

F2x = k q2 ( q3 - q4 ) sin \theta/ d2

= 9 x 109 x 5.9 x 10-6 x sin60 x ( 0 ) x 10-6 / 0.0932

F2x = 0 N

e )

F1x = k q1 ( q2 - q3 - q4 ) cos \theta/ d2  

= 9 x 109 x 5.8 x 10-6 x cos60 x (5.9 - 3.4 - 3.4 ) x 10-6 / 0.0932

F1x = - 4.7 N